# On dense subalgebras of the singular ideal in groupoid C*-algebras

### v1

@@ 1.2. Hausdorff cover
Elements of $\tilde{G}$ can be described explicitly as follows. A non-empty subset $\bm{g}\subseteq G$ lies in $\tilde{G}$ if and only if there exist a net $(g_{i})$ in $G$ such that $\bm{g}$ is the set of limit points of $(g_{i})$, and every limit point of a subnet of $(g_{i})$ (cluster point) is a limit point of $(g_{i})$. Moreover, given a net $(g_{i})$ in $G$, the image net $\iota(g_{i})$ converges to $\bm{g}\in\tilde{G}$ in the topology of $\tilde{G}$ if and only if the conditions above are satisfied.
Define $\tilde{G}^{(0)}\coloneq\{\bm{g}\in\tilde{G}\colon\bm{g}\cap G^{(0)}\neq\emptyset\}$. This set will be the unit space of $\tilde{G}$ with the structure described below. The canonical inclusion map $\iota\colon G\hookrightarrow\tilde{G}$ restricts to an inclusion $G^{(0)}\hookrightarrow\tilde{G}^{(0)}$, and the image $\iota(G^{(0)})$ is dense in $\tilde{G}^{(0)}$. For each $\bm{x}\in\tilde{G}^{(0)}$ there is a unique element $\pi(\bm{x})\in\bm{x}\cap G^{(0)}$, and the map $\pi\colon\tilde{G}^{(0)}\to G^{(0)}$ is a continuous surjection. Since $\iota(G^{(0)})$ is dense in $\tilde{G}^{(0)}$, continuity of the groupoid operations on $G$ implies that any element $\bm{x}\in\tilde{G}^{(0)}$ is a subgroup of the isotropy group $G_{\pi(\bm{x})}^{\pi(\bm{x})}$. We will often write $\bm{x}=X$ to distinguish elements $x\in G^{(0)}$ from $\bm{x}\in\tilde{G}^{(0)}$.
We now describe the groupoid operations on $\tilde{G}$. Note that $r(g_{1})=r(g_{2})$ and $s(g_{1})=s(g_{2})$ whenever $g_{1},g_{2}\in\bm{g}$ and $\bm{g}\in\tilde{G}$. The fact that elements of $\tilde{G}^{(0)}$ are subgroups implies, for every $\bm{g}\in\tilde{G}$, there are unique elements $\bm{x},\bm{y}\in\tilde{G}^{(0)}$ such that $\bm{g}=g\cdot\bm{x}=\bm{y}\cdot g$ for any $g\in\bm{g}$. Let $s(\bm{g})=\bm{x}$ and $r(\bm{g})=\bm{y}$. Define the inverse $\bm{g}^{-1}\coloneq\{g^{-1}\colon g\in\bm{g}\}$ and the product $\bm{g}\bm{h}\coloneq\{gh\colon g\in\bm{g},h\in\bm{h}\}$ whenever $s(\bm{g})=r(\bm{h})$. Note also that
$\bm{g}\bm{h}=\{gh\colon h\in\bm{h}\}\text{ for any }g\in\bm{g}.$ (1.1)
With this structure, $\tilde{G}$ becomes a Hausdorff étale groupoid.
@@
$\mathfrak{i}(f)(\bm{g})=\sum_{g\in\bm{g}}f(g)$ (1.2)
for all $\bm{g}\in\tilde{G}$ ([4, Lemma 3.2]). The embedding $Cc(G~)\mathfrak{i}\colon\mathscr$$\hookrightarrow C_{c}(\tilde{G})$ is a -homomorphism, and induces canonical embeddings $Cr(G~)\mathfrak{i}_{r}\coloneq C^{*}_{r}(\mathfrak{i})\colon$ $\hookrightarrow C^{*}_{r}(\tilde{G})$ and $C(G~)\mathfrak{i}\coloneq C^{*}(\mathfrak{i})\colon$ $\hookrightarrow C^{*}(\tilde{G})$ ([4, Lemma 3.8 & Corollary 6.8]).
Now assume that the groupoid $G$ is covered by countably many open bisections. We say a unit $x\in G^{(0)}$ is Hausdorff if $\pi^{-1}(x)=\{\iota(x)\}$, and we let $C\subseteq G^{(0)}$ denote the set of Hausdorff units. By [19, Lemma 7.15], $C$ is dense in $G^{(0)}$ (the Hausdorff points complement the so-called dangerous points of [19]). Define $\tilde{G}^{(0)}_{\rm{ess}}\coloneq\overline{\iota(C)}$. This is an invariant subset of the unit space $\tilde{G}^{(0)}$. We define the essential Hausdorff cover $\tilde{G}_{\rm{ess}}\coloneq\tilde{G}|_{\tilde{G}^{(0)}_{\rm{ess}}}=\{\bm{g}\in\tilde{G}\colon s(\bm{g})\in\tilde{G}^{(0)}_{\rm{ess}}\}$ ([4, Definition 4.13]).
@@ 1.3. Isotropy fibres of ideals
Our attention now turns to restriction maps. Let $G$ be an étale groupoid. Given a unit $x\in G^{(0)}$, consider the restriction map $\eta_{x}\colon\mathscr{C}_{c}(G)\to\mathbb{C}[G^{x}_{x}]$, where $\mathbb{C}[G^{x}_{x}]$ denotes the group algebra of the isotropy group $G^{x}_{x}$ (thought of as finitely supported functions $G^{x}_{x}\to\mathbb{C}$). The map $\eta_{x}$ extends to a completely positive contraction $C^{*}(G)\to C^{*}(G^{x}_{x})$, which we will also denote by $\eta_{x}$. This can be seen using induced representations of the isotropy group (see [7, Lemma 1.2] for example). In fact, any $C^{*}$-norm on $\mathscr{C}_{c}(G)$ dominating the reduced norm will define a $C^{*}$-norm on the group algebra $\mathbb{C}[G^{x}_{x}]$ dominating the reduced norm, and the restriction map $\mathscr{C}_{c}(G)\to\mathbb{C}[G^{x}_{x}]$ will always extend to a completely positive contraction between the associated $C^{*}$-algebras. See [8, Section 1.3] for a discussion in this direction. For a (closed) ideal $I\trianglelefteq C^{*}(G)$, the image $\eta_{x}(I)$ is also a (closed) ideal in $C^{*}(G^{x}_{x})$ ([8, Lemma 2.1]). We write $I_{x}\coloneq\eta_{x}(I)$ and, following [8], call $I_{x}$ the isotropy fibre of $I$ at $x$.
For $g\in G^{x}_{x}$, denote by $\delta_{g}\in\mathbb{C}[G^{x}_{x}]$ the function equal to one at $g$ and zero everywhere else. Each $g\in G$ induces an isomorphism of group algebras $\text{Ad}_{g}\colon\mathbb{C}\big[G^{s(g)}_{s(g)}\big]\to\mathbb{C}\big[G^{r(g)}_{r(g)}\big]$, $\delta_{h}\mapsto\delta_{ghg^{-1}}$ and this maps extends to an isomorphism of $C^{*}$-algebras $\Psi_{g}\colon C^{*}\big(G^{s(g)}_{s(g)}\big)\to C^{*}\big(G^{r(g)}_{r(g)}\big)$. A family $(K_{x})_{x\in G^{(0)}}$ of ideals satisfying $K_{x}\trianglelefteq C^{*}(G_{x}^{x})$ for each $x\in G^{(0)}$ is said to be invariant if $\Psi_{g}(K_{s(g)})=K_{r(g)}$ for all $g\in G$. Given an ideal $I\trianglelefteq C^{*}(G)$, its isotropy fibres form an invariant family of ideals. Going in the other direction, if $\mathcal{K}=(K_{x})_{x\in G^{(0)}}$ is an invariant family of ideals, we define
@@
By [8, Corollary 2.9], $I(\mathcal{K})$ is an ideal in $C^{*}(G)$. We isolate the following observation from [8].
@@
Let $\mathcal{K}=(K_{x})_{x\in G^{(0)}}$ be an invariant family of ideals, and let $A\subseteq C^{*}(G)$ be a sub-$C^{*}$-algebra. Then $A\subseteq I(\mathcal{K})$ if and only if $\eta_{x}(A)\subseteq K_{x}$ for all $x\in G^{(0)}$.
@@
Assume that $A$ satisfies $A\subseteq I(\mathcal{K})$. Fix $x\in G^{(0)}$. We have that $\eta_{x}(a)\in K_{x}$ for all positive elements $a\in A^{+}$. Since $A$ is spanned by its positive elements, it follows that $\eta_{x}(a)\in K_{x}$ for all $a\in A$. Conversely, assume that $\eta_{x}(A)\subseteq K_{x}$ for all $x\in G^{(0)}$. For any $a\in A$ we have $\eta_{x}(a^{*}a)\in K_{x}$ for all $x\in G^{(0)}$, implying that $a\in I(\mathcal{K})$ as desired. ∎
@@ 2. Ideals are determined by their isotropy fibres
Let $X\in\tilde{G}^{(0)}$ and write $x\coloneq\pi(X)\in G^{(0)}$. By definition, $X$ is equal to a subgroup of the isotropy group $G^{x}_{x}$. It is easy to see that the isotropy group $\tilde{G}^{X}_{X}\subseteq\tilde{G}$ is equal to the quotient $N_{X}/X$, where $N_{X}=\{g\in G^{x}_{x}:gXg^{-1}=X\}$ is the normaliser of $X$ in $G^{x}_{x}$. Denote by $E_{N_{X}}\colon C^{*}(G^{x}_{x})\to C^{*}(N_{X})$ the canonical conditional expectation and $Q_{X}\colon C^{*}(N_{X})\to C^{*}(N_{X}/X)$ the -homomorphism induced by the quotient map $N_{X}\to N_{X}/X$.
@@
Equality is verified readily on the dense *-subalgebra $\mathscr{C}_{c}(G)$, and the lemma follows immediately. ∎
@@
$\Psi_{\bm{g}}(K_{s(\bm{g})})$ $=\overline{\Psi_{\bm{g}}\big(Q_{s(\bm{g})}\circ E_{N_{s(\bm{g})}}(I_{s(g_{0})})\big)}=\overline{Q_{r(\bm{g})}\circ E_{N_{r(\bm{g})}}\big(\Psi_{g_{0}}(I_{s(g_{0})})\big)}$
$=\overline{Q_{r(\bm{g})}\circ E_{N_{r(\bm{g})}}(I_{r(g_{0})})}=K_{r(\bm{g})}.$
@@
By Lemma 2.3, we have $K_{X}=\overline{\eta_{X}\big(\mathfrak{i}(I)\big)}$, and therefore $K_{X}\subseteq\tilde{I}_{X}$ for all $X\in\tilde{G}^{(0)}$. By Lemma 1.3, this equality also implies that $\mathfrak{i}(I)\subseteq I(\mathcal{K})$. Hence $\tilde{I}\subseteq I(\mathcal{K})$, and another application of Lemma 1.3 gives $\tilde{I}_{X}\subseteq K_{X}$ for all $X\in\tilde{G}^{(0)}$, as desired. ∎
We now prove that ideals in $C^{*}(G)$ are determined by their isotropy fibres whenever $G$ is an amenable second-countable étale groupoid. This generalises [8, Theorem 2.10] to the setting of non-Hausdorff groupoids. Our proof utilises this theorem, which in turn uses known cases of the Effros-Hahn conjecture [17]. Therefore it is necessary for us to only consider amenable second-countable groupoids. The theorem does not hold in the non-amenable setting - counterexamples are provided in [8, Examples 2.11 & 2.12].
@@
The forwards implication is trivial. Assume $I_{x}=K_{x}$ for all $x\in G^{(0)}$. Proposition 2.4 implies that $\tilde{I}_{X}=\tilde{K}_{X}$ for all $X\in\tilde{G}^{(0)}$. Since $G$ is amenable and second-countable, so is $\tilde{G}$ by [4, Theorem 6.5]. Moreover, the groupoid $\tilde{G}$ is Hausdorff, and therefore [8, Theorem 2.10] ensures that $\tilde{I}=\tilde{K}$. Finally, Proposition 2.2 gives $I=\mathfrak{i}^{-1}(\tilde{I})=\mathfrak{i}^{-1}(\tilde{K})=K$. ∎
2.1. Characterisation of density
In this subsection, we assume that $G$ is an amenable second-countable (not necessarily Hausdorff) étale groupoid, and therefore $C^{*}(G)=C^{*}_{r}(G)$ by [4, Corollary 6.10]. We denote by $J$ the singular ideal in $C^{*}_{r}(G)$. Theorem 2.5 yields the following characterisation for the density of $J\cap\mathscr{C}_{c}(G)$ in $J$.

### v2

@@ 1.2. Hausdorff cover
Elements of $\tilde{G}$ can be described explicitly as follows. A non-empty subset $\bm{g}\subseteq G$ belongs to $\tilde{G}$ if and only if there exists a net $(g_{\alpha})$ in $G$ whose set of limit points is precisely $\bm{g}$, and such that every subnet of $(g_{\alpha})$ has its limit points contained in (and hence equal to) $\bm{g}$. Moreover, given a net $(g_{\alpha})$ in $G$, the image net $\iota(g_{\alpha})$ converges to $\bm{g}\in\tilde{G}$ in the Fell topology if and only if the conditions above are satisfied. We will use the following fact several times: any convergent net $(g_{\alpha})$ in $G$ has a subnet $(g_{\beta})$ such that $\iota(g_{\beta})$ is convergent in $\tilde{G}$.
Define $\tilde{G}^{(0)}\coloneq\{\bm{g}\in\tilde{G}\colon\bm{g}\cap G^{(0)}\neq\emptyset\}$. This set will be the unit space of $\tilde{G}$ with the groupoid structure described below. The canonical inclusion map $\iota\colon G\hookrightarrow\tilde{G}$ restricts to an inclusion $G^{(0)}\hookrightarrow\tilde{G}^{(0)}$, and the image $\iota(G^{(0)})$ is dense in $\tilde{G}^{(0)}$. In particular, any element $\bm{x}\in\tilde{G}^{(0)}$ is contained in the closure $\overline{G^{(0)}}$. For each $\bm{x}\in\tilde{G}^{(0)}$ there is a unique element $\pi(\bm{x})\in\bm{x}\cap G^{(0)}$, and the map $\pi\colon\tilde{G}^{(0)}\to G^{(0)}$ is a proper continuous surjection (since $\iota(G^{(0)})$ is dense in $\tilde{G}^{(0)}$). Continuity of the groupoid operations on $G$ implies that any element $\bm{x}\in\tilde{G}^{(0)}$ is a subgroup of the isotropy group $G_{\pi(\bm{x})}^{\pi(\bm{x})}$. We will often write $\bm{x}=X$ to distinguish elements $x\in G^{(0)}$ from $\bm{x}\in\tilde{G}^{(0)}$.
We now describe the groupoid operations on $\tilde{G}$. With these operations, $\tilde{G}$ becomes a Hausdorff étale groupoid. Note that $r(g_{1})=r(g_{2})$ and $s(g_{1})=s(g_{2})$ whenever $g_{1},g_{2}\in\bm{g}$ and $\bm{g}\in\tilde{G}$. Since elements of $\tilde{G}^{(0)}$ are subgroups of isotropy groups, we have that for any $\bm{g}\in\tilde{G}$, there are unique elements $X,Y\in\tilde{G}^{(0)}$ such that $\bm{g}=g\cdot X=Y\cdot g$ for all $g\in\bm{g}$. Let $s(\bm{g})\coloneq X$ and $r(\bm{g})\coloneq Y$. Explicitly, we have
(1.1) $s(\bm{g})=g^{-1}\cdot\bm{g}\quad\quad\text{and}\quad\quad r(\bm{g})=\bm{g}\cdot g^{-1}\quad\quad\text{for any }g\in\bm{g}.$
Define the inverse $\bm{g}^{-1}\coloneq\{g^{-1}\colon g\in\bm{g}\}$ and the product $\bm{g}\bm{h}\coloneq\{gh\colon g\in\bm{g},h\in\bm{h}\}$ whenever $s(\bm{g})=r(\bm{h})$. Note that
@@
[…]
(1.3) $\mathfrak{i}(f)(\bm{g})=\sum_{g\in\bm{g}}f(g)$
for all $\bm{g}\in\tilde{G}$ ([5, Lemma 3.2]). The embedding $Cc(G~)\mathfrak{i}\colon\mathscr$$\hookrightarrow C_{c}(\tilde{G})$ is a -homomorphism, and induces canonical embeddings $Cr(G~)\mathfrak{i}_{r}\coloneq C^{*}_{r}(\mathfrak{i})\colon$ $\hookrightarrow C^{*}_{r}(\tilde{G})$ and $C(G~)\mathfrak{i}\coloneq C^{*}(\mathfrak{i})\colon$ $\hookrightarrow C^{*}(\tilde{G})$ ([5, Lemma 3.8 & Corollary 6.8]).
Now assume that the groupoid $G$ is covered by countably many open bisections. Observe that a unit $x\in G^{(0)}$ is Hausdorff if and only if the inclusion $\iota\colon G^{(0)}\to\tilde{G}^{(0)}$ is continuous at $x$, if and only if $\pi^{-1}\big(\{x\}\big)=\{\iota(x)\}$. Define $\tilde{G}^{(0)}_{\rm{ess}}\coloneq\overline{\iota(C)}$, where $C\subseteq G^{(0)}$ is the set of Hausdorff units. This is an invariant subset of the unit space $\tilde{G}^{(0)}$ of the Hausdorff cover. We define the essential Hausdorff cover $\tilde{G}_{\rm{ess}}\coloneq\tilde{G}|_{\tilde{G}^{(0)}_{\rm{ess}}}=\{\bm{g}\in\tilde{G}\colon s(\bm{g})\in\tilde{G}^{(0)}_{\rm{ess}}\}$ ([5, Definition 4.13]).
Definition 1.3 ([19]).
Let $G$ be an étale groupoid that is covered by countably many open bisections. For each $x\in G^{(0)}$ define $\mathcal{X}(x)\coloneq\pi^{-1}\big(\{x\}\big)\cap\tilde{G}^{(0)}_{{\rm{ess}}}$.
By [19], $\mathcal{X}(x)$ is a conjugation-invariant collection of subgroups of $G^{x}_{x}$ that is closed in the Chabauty topology (see Section 3). A subgroup $X\subseteq G^{x}_{x}$ belongs to $\mathcal{X}(x)$ if and only if there exists a net $(x_{\alpha})\subset C$ of (Hausdorff) units whose set of limit points is precisely $X$, and such that every subnet of $(x_{\alpha})$ has its limit points contained in $X$.
$\bullet$
@@ 1.3. Isotropy fibres of ideals
Our attention now turns to restriction maps. Let $G$ be an étale groupoid. Given a unit $x\in G^{(0)}$, consider the restriction map $\eta_{x}\colon\mathscr{C}_{c}(G)\to\mathbb{C}[G^{x}_{x}]$, where $\mathbb{C}[G^{x}_{x}]$ denotes the group algebra of the isotropy group $G^{x}_{x}$ (thought of as finitely supported functions $G^{x}_{x}\to\mathbb{C}$). The map $\eta_{x}$ extends to a completely positive contraction $C^{*}(G)\to C^{*}(G^{x}_{x})$, which we will also denote by $\eta_{x}$. This can be seen using induced representations of the isotropy group (see [8, Lemma 1.2] for example). For a (closed) ideal $I\trianglelefteq C^{*}(G)$, the image $\eta_{x}(I)$ is also a (closed) ideal in $C^{*}(G^{x}_{x})$ ([9, Lemma 2.1]). We write $I_{x}\coloneq\eta_{x}(I)$ and, following [9], call $I_{x}$ the isotropy fibre of $I$ at $x$.
For $g\in G^{x}_{x}$, denote by $\delta_{g}\in\mathbb{C}[G^{x}_{x}]$ the function equal to 1 at $g$ and zero everywhere else. Each $g\in G$ induces an isomorphism of group algebras $\text{Ad}_{g}\colon\mathbb{C}\big[G^{s(g)}_{s(g)}\big]\to\mathbb{C}\big[G^{r(g)}_{r(g)}\big]$; $\delta_{h}\mapsto\delta_{ghg^{-1}}$ and this map extends to an isomorphism of $C^{*}$-algebras $\Psi_{g}\colon C^{*}\big(G^{s(g)}_{s(g)}\big)\to C^{*}\big(G^{r(g)}_{r(g)}\big)$. A family $(K_{x})_{x\in G^{(0)}}$ of ideals satisfying $K_{x}\trianglelefteq C^{*}(G_{x}^{x})$ for each $x\in G^{(0)}$ is said to be invariant if $\Psi_{g}(K_{s(g)})=K_{r(g)}$ for all $g\in G$. Given an ideal $I\trianglelefteq C^{*}(G)$, its isotropy fibres form an invariant family of ideals. Going in the other direction, if $\mathcal{K}=(K_{x})_{x\in G^{(0)}}$ is an invariant family of ideals, we define
@@
By [9, Corollary 2.9], $I(\mathcal{K})$ is an ideal in $C^{*}(G)$. We isolate the following observation from [9].
@@
Let $\mathcal{K}=(K_{x})_{x\in G^{(0)}}$ be an invariant family of ideals, and let $A\subseteq C^{*}(G)$ be a $C^{*}$-subalgebra. Then $A\subseteq I(\mathcal{K})$ if and only if $\eta_{x}(A)\subseteq K_{x}$ for all $x\in G^{(0)}$.
@@
Assume that $A$ satisfies $A\subseteq I(\mathcal{K})$. Fix $x\in G^{(0)}$. We have $\eta_{x}(a)\in K_{x}$ for all positive elements $a\in A^{+}$. Since $A$ is spanned by its positive elements, it follows that $\eta_{x}(a)\in K_{x}$ for all $a\in A$. Conversely, assume that $\eta_{x}(A)\subseteq K_{x}$ for all $x\in G^{(0)}$. For any $a\in A$ we have $\eta_{x}(a^{*}a)\in K_{x}$ for all $x\in G^{(0)}$, and this implies that $a\in I(\mathcal{K})$ as desired. ∎
@@ 2. Ideals are determined by their isotropy fibres
Let $X\in\tilde{G}^{(0)}$ and write $x\coloneq\pi(X)\in G^{(0)}$. By definition, $X$ is equal to a subgroup of the isotropy group $G^{x}_{x}$. By (1.1) it is easily seen that the isotropy group $\tilde{G}^{X}_{X}\subseteq\tilde{G}$ is equal to the quotient $N_{X}/X$, where $N_{X}=\{g\in G^{x}_{x}\colon gXg^{-1}=X\}$ is the normaliser of $X$ in $G^{x}_{x}$. Denote by $E_{N_{X}}\colon C^{*}(G^{x}_{x})\to C^{*}(N_{X})$ the canonical conditional expectation and by $Q_{X}\colon C^{*}(N_{X})\to C^{*}(N_{X}/X)$ the -homomorphism induced by the quotient map $N_{X}\to N_{X}/X$.
@@
Equality is readily verified on the dense -subalgebra $\mathscr{C}_{c}(G)$. The lemma then follows by continuity. ∎
@@
$\Psi_{\bm{g}}\big(K_{s(\bm{g})}\big)$ $=\overline{\Psi_{\bm{g}}\Big(Q_{s(\bm{g})}\circ E_{N_{s(\bm{g})}}\big(I_{s(g_{0})}\big)\Big)}=\overline{Q_{r(\bm{g})}\circ E_{N_{r(\bm{g})}}\Big(\Psi_{g_{0}}\big(I_{s(g_{0})}\big)\Big)}$
$=\overline{Q_{r(\bm{g})}\circ E_{N_{r(\bm{g})}}\big(I_{r(g_{0})}\big)}=K_{r(\bm{g})}.$
@@
By Lemma 2.3, we have $K_{X}=\overline{\eta_{X}\big(\mathfrak{i}(I)\big)}$, and therefore $K_{X}\subseteq\tilde{I}_{X}$ for all $X\in\tilde{G}^{(0)}$. By Lemma 1.4, this equality also implies that $\mathfrak{i}(I)\subseteq I(\mathcal{K})$, and hence $\tilde{I}\subseteq I(\mathcal{K})$. Another application of Lemma 1.4 gives $\tilde{I}_{X}\subseteq K_{X}$ for all $X\in\tilde{G}^{(0)}$, as desired. ∎
We now prove that ideals in $C^{*}(G)$ are determined by their isotropy fibres whenever $G$ is an amenable second-countable étale groupoid. This generalises [9, Theorem 2.10] to the setting of non-Hausdorff groupoids. Our proof utilises this theorem, which in turn uses known cases of the Effros-Hahn conjecture [20]. Therefore, it is necessary for us to only consider amenable second-countable groupoids. The theorem does not hold in the non-amenable setting – counterexamples are provided in [9, Examples 2.11 & 2.12].
@@
The forwards implication is trivial. Assume $I_{x}=K_{x}$ for all $x\in G^{(0)}$. Proposition 2.4 implies that $\tilde{I}_{X}=\tilde{K}_{X}$ for all $X\in\tilde{G}^{(0)}$. Since $G$ is amenable and second-countable, so is $\tilde{G}$ by [5, Theorem 6.5]. Moreover, the groupoid $\tilde{G}$ is Hausdorff, and therefore [9, Theorem 2.10] ensures that $\tilde{I}=\tilde{K}$. Finally, Proposition 2.2 gives $I=\mathfrak{i}^{-1}(\tilde{I})=\mathfrak{i}^{-1}(\tilde{K})=K$. ∎
3. Characterisation of density
In this section, we assume that $G$ is an amenable second-countable (not necessarily Hausdorff) étale groupoid, and therefore $C^{*}(G)=C^{*}_{r}(G)$ by [5, Corollary 6.10]. We denote by $J$ the singular ideal in $C^{*}_{r}(G)$. Theorem 2.5 yields the following characterisation for the density of $J\cap\mathscr{C}_{c}(G)$ in $J$.